OpenStudy (anonymous):

What is the equation of a line that passes through the point (– 3,–11) and is parallel to the line whose equation is 2x – y = 4?

OpenStudy (anonymous):

slope of \[2x-y=4\] is 2

OpenStudy (anonymous):

so your job is to find the equation of the line with slope 2 through (-3,-11) use \[y-y_1=m(x-x_1)\] with \[m=2, x_1=-3,y_1=-11\] get \[y+11=2(x+3)\]

OpenStudy (anonymous):

y=2x-5